OpenStudy (anonymous):

A dentist selected 16 patients at random from a population of more than 1000 patients and asked them how many times the patient flosses per week. The dentist made a line plot of the results. What is the estimated mode of the number of times all patients floss per week?

OpenStudy (anonymous):

A. 5 B. 5.5 C. 6 D. 7

OpenStudy (anonymous):

So we have 7 five times , 6 three times, 5 four times and 1 once and 2 one time we don't need the zeros

OpenStudy (anonymous):

7 *5 = 35 6*3= 18 5*4= 20 2*1= 2 1*1=1 If u add all it will be 76 then divide by 16

OpenStudy (anonymous):

What is 76 divided by 16?

OpenStudy (anonymous):

4.75

OpenStudy (anonymous):

Hold on I thnk I typed that out wrong but the answer is 7

OpenStudy (anonymous):

Okey...Thank you C:

OpenStudy (anonymous):

ok the mode is the most so basically the one with the most x's on the plot should be your answer i just tried to do it mathematically and that didnt work out so XD but 7 is your answer